Suppose that a change in the price of bread from R10 to R12.50 leads to a change in the quantity demanded of bread from 65 units to 90 .use the arc elasticity method to calculate and interpret the cross price elasticity of demand for the bread


1
Expert's answer
2020-10-13T10:45:36-0400

Using the arc elasticity method the elasticity is computed using the following formula:

"\text{Arc elasticity}=\dfrac{\dfrac{Q_2-Q_1}{\frac{Q1+Q_2}{2}}}{{\dfrac{P_2-P_1}{\frac{P_1+P_2}{2}}}}"



"\text{Arc elasticity}=\dfrac{\dfrac{90-65}{\frac{65+90}{2}}}{{\dfrac{12.5-10}{\frac{10+12.5}{2}}}}"


"\text{Arc elasticity}=\blue{1.45}"


The cross price elasticity of demand will be:

"\text{Cross price elasticity of damand}=\dfrac{\delta Q}{\delta P} \times \dfrac{P}{Q}"


"\text{Cross price elasticity of damand}=\dfrac{90-65}{12.5-10} \times \dfrac{12.5}{90}"


"\text{Cross price elasticity of damand}=\blue{1.33}"